Determine the average rate of change of B from t = 0s to t =392s.

A⟶2B

Time(s) ---- Concentration of A (M)
0 0.720
196 0.430
392 0.140
1
Expert's answer
2020-03-02T06:57:02-0500

1) Find the average rate of disappearance of A

"Rate 1 (A) = -\frac{\Delta C}{\Delta t}= -\frac{0.430-0.720}{196-0}=\frac{0.29}{196}=1.48\times 10^{-3} \frac{mol}{L\times s}"


"Rate2(A)= -\frac{\Delta c}{\Delta t}=\frac{0.140-0.430}{392-196}= \frac{0.29}{196}=1.48\times10^{-3}\frac{mol}{L\times s}"


Average rate of A = "1.48\times10^{-3}\frac{mol}{L\times s}"


Since A-> 2B, then to find the average rate of appearance of B we need to multiply the average rate of disappearance of A by 2:

Rate of B ="1.48\times 10^{-3}\times 2= 2.96\times 10^{-3}\frac{mol}{L\times s}"